OpenStudy (anonymous):

Given h = -2 and k = 4, what is the equation of the graph if the parent function is y equals the square root of x? Describe the transformation.

OpenStudy (anonymous):

@dauspex

OpenStudy (anonymous):

ERR Start with y = sqrt (x) as h = -2, sub y with y - h, ie y + 2 as k = 4, sub x with x -k , ie x -4 Thus you'll get y+2 = sqrt (x-4)

OpenStudy (anonymous):

ERR Start with y = sqrt (x) as h = -2, sub y with y - h, ie y + 2 ===> graph is shifted two spaces up as k = 4, sub x with x -k , ie x -4 ===> graph is shift 4 spaces right Thus you'll get y+2 = sqrt (x-4)

OpenStudy (anonymous):

thanks mate! @dauspex